A 100 N force that makes an angle of 30with the horizontal is used to move a 10 kg crate across a smooth floor. The crate moves through a distance of 500 cm. Find the work done by the force.


1
Expert's answer
2022-04-17T17:08:05-0400

The work done by the force

"W=Fd\cos\theta\\
=100\:\rm N*5\: m*\cos30^\circ=433\: J"